# Word length predicts word order: “Min-max”-ing drives language evolution

### v1

@@ Abstract
Current theories of language propose an innate (Baker 2001; Chomsky 1981) or a functional (Greenberg 1963; Dryer 2007; Hawkins 2014) origin for the surface structures (i.e. word order) that we observe in languages of the world, while evolutionary modeling (Dunn et al. 2011) suggests that descent is the primary factor influencing such patterns. Although there are hypotheses for word order change from both innate and usage-based perspectives for specific languages and families, there are key disagreements between the two major proposals for mechanisms that drive the evolution of language more broadly (Wasow 2002; Levy 2008). This paper proposes a universal underlying mechanism for word order change based on a large tagged parallel dataset of over 1,500 languages representing 133 language families and 111 isolates. Results indicate that word class length is significantly correlated with word order crosslinguistically, but not in a straightforward manner, partially supporting opposing theories of processing, while at the same time predicting historical word order change in two different phylogenetic lines and explaining more variance than descent or language area in regression models. Such findings suggest an integrated “Min-Max” theory of language evolution driven by competing pressures of processing and information structure, aligning with recent efficiency-oriented (Levshina 2023) and information-theoretic proposals (Zaslavsky 2020; Tucker et al. 2025).
@@ 1 Introduction
Theories of language should describe not only the current state of languages, but also explain how they have developed (their history) and predict how they will change (their future). To date, however, existing theories have primarily focused on describing the present state of language, proposing that the current diversity in surface structures or patterns that we observe in languages of the world, such as for word order, have either an innate (Baker 2001; Chomsky 1981) or functional (Greenberg 1963; Dryer 2007; Hawkins 2014) source. More recent diachronic work has tried to answer the second question on how languages develop. From an “innateness” perspective, surface word orders are seen to be derived from an underlying (universally shared) structure, whereby parameters for surface realizations change over time (Carnie & Guilfoyle 2000; Crisma & Longobardi 2009; Martins & Cardoso 2018), while from a “functional” perspective surface word orders are the result of cumulative changes in linked structures (sometimes called ‘constructions’) that compete for adoption in individual languages (see Gerritsen & Stein 1992; Leino & Kuningas 2006; Barðdal & Gildea 2016).
Importantly, none of these perspectives are particularly explanatory of drivers in language evolution that affect all languages similarly. The closest observations in this regard come from evolutionary modeling, where descent from a common ancestor was found to be the most explanatory factor in predicting word order (Dunn et al. 2011; though see Jäger & Wahle 2021; Naranjo & Becker 2022 for alternate perspectives), and language processing, where word length (Piantadosi et al. 2011; Wasow 2022) and surprisal (Levy 2008) have been shown to impact the realization of word order. There are also more general claims that “efficiency” (variously defined) explains grammatical (Hawkins 2014; Levshina 2023) or semantic (Zaslavsky 2020) change. However, the fact that languages tend to display word orders like their family members does not explain the variation in word orders that we find within families, and perspectives on processing make very different predictions: the literature on word length predicts that short/light words will occur first in a sentence, while the literature on surprisal predicts that long/heavy words will occur first. Further, the literature on efficiency does not make actual testable cross-linguistic predictions about basic word order.
There are also other issues to consider, based on observation of the languages for which we have data. Typological surveys (Dryer & Haspelmath 2013; Skirgård et al. 2023; Bickel et al. 2023) reveal that over 80% of the sampled languages have a Subject-initial basic word order, while only about 10% have a Verb-initial order. Further, there are various correlations to be found between the orders of clausal elements, such that basic word order is related to how other clause types are ordered. While all languages allow variation in word order to different degrees, the word order most often found in declarative sentences can predict (to some degree) the ordering of other grammatical elements in a language (Dryer 1992, 2011).
It is not completely clear why one basic constituent order correlates better with one particular order of (e.g.) noun phrase elements than another, nor why some basic word orders are less well-represented crosslinguistically. This raises the question whether there are underlying processes that drive these observed tendencies. Some explanation clearly comes from descent, i.e. that languages tend to resemble their ancestral forms. However, the word order changes that languages undergo over time suggests other pressures in terms of how complex linguistic elements are and how these constituents are processed in the brain (Maurits & Griffiths 2014; Schouwstra & de Swart 2014, see also Piantadosi et al. 2011; Bickel et al. 2015; Ding et al. 2017; Fernandez 2016; Jaeger & Tily 2011; Tily 2010; Sinnemäki 2014).
Answers to these questions have been elusive in part due to a lack of data that would facilitate crosslinguistic investigations. This is because such investigations require data that is annotated for various word types or dependencies, which is unavailable for the majority of the world’s languages. The largest sample of languages used to investigate these questions looked at annotated corpora in 80 languages (Hahn & Xu 2022), the majority of which are from the Indo-European language family. Such data risks biasing findings toward a particular group of languages (Blasi et al. 2022), in much the same way that other scientific findings have historically been biased toward WEIRD demographics (Henrich et al. 2010), leading to recent calls for more expansive datasets to examine linguistic questions (Benítez-Burraco 2025; Levshina et al. 2023).
The current study addresses this by developing a large corpus of texts tagged for word class. With data sourced from the Parallel Bible Corpus (PBC; Mayer & Cysouw 2014), machine learning techniques allow for automatic tagging [^1] Automated part of speech tagging was validated via comparison with existing taggers for high-resource languages (SpaCy, 16 languages - avg nouns: 0.83, verbs: 0.74; Trankit, 31 languages - avg nouns: 0.83, verbs: 0.76) and gold-standard tagged data in the Universal Dependencies Treebanks (UDT; Zeman et al. 2024; 106 languages; https://universaldependencies.org). See Ring (2025) for more details. of word classes in 1,597 languages (roughly 20% of the world’s languages) that represent 133 families and 111 isolates.222‘Isolates’ being defined here as languages in the corpus that share no known family membership with other languages in the corpus. Individual corpora in the majority of these languages (over 1,500) contain more than 1,800 parallel sentences, and the smallest corpus contains more than 700 sentences. [^3] For clarity, the UDT has 160+ hand-annotated languages available for crosslinguistic investigation, but these are not parallel texts and the number of sentences in each language is variable, with some having fewer than 200 sentences in the v2.14 release. The ODIN project (and its re-packaged form via LECS: https://huggingface.co/datasets/lecslab/glosslm-corpus) has interlinear glossed text (IGT) from over 1,800 languages, but the majority of these are single-word/line examples, and only 300 of the languages in their dataset have more than 100 sentences. Additional cross-linguistic investigations using tagged corpora include Jing et al. (2021: 71 languages) and Hahn et al. (2020: 51 languages). This dwarfs both the linguistic data and the language diversity available from other sources and means that crosslinguistic investigations using this dataset can be confident that they are comparing similar contexts. The dataset (the tagged PBC) and explanations regarding its development, along with Python code for replicating the results in this paper, can be found at the linked repository [^4] https://github.com/lingdoc/taggedPBC and is described at greater length in Ring (2025).
This dataset can be used to investigate the crosslinguistic questions highlighted above regarding whether there is some pressure (beyond descent) that explains the evolution of word order in languages of the world. If there is indeed a mechanism that underlies the observed surface structures in present-day languages, this should also be one that reveals historical processes and allows us to predict future developments. Specifically, by deriving a novel corpus-based measure of word order (the “N1 ratio”) [^5] The N1 ratio was validated via comparison with existing classifications of languages in the tagged PBC and three typological databases: the World Atlas of Language Structures (WALS; Dryer & Haspelmath 2013; https://wals.info/), the AUTOTYP Database (Bickel et al. 2023; https://github.com/autotyp/autotyp-data), and Grambank (Skirgård et al. 2023; https://grambank.clld.org/) where the N1 ratio significantly differentiated (p ¡ 0.001) between SV/VS/free classifications in all three databases. See Ring (2025) for more details. from this dataset, the current paper investigates the relationship between basic word order and the lengths of particular word classes (nouns/arguments and verbs/predicates) crosslinguistically. Results show that the lengths of these word classes is significantly related to word order, that they can predict historical word order change in two different lineages, and that they account for more variance than family membership (descent) or areal patterns (geography) when examined using hierarchical linear regressions. Such findings suggest the possibility for a unified (“Min-Max”) theory of language evolution that can explain both observed surface structures in languages of the world as well as how such patterns have developed diachronically.
2 Complexity, processing, and word order
Linguistic complexity can take many forms, with many considerations regarding its measurement (see Jaeger & Tily 2011). As noted by Joseph and Newmayer (2012: 341):
“how linguistic complexity is conceived of, and how it might be identified and measured, has itself never been agreed upon by linguists. It cannot be neatly separated from theoretical considerations of where language structure ends and language processing (or use) begins, or where variation (whether social or stylistic) is located.”
Some current efforts to measure linguistic complexity focus on quantifying the overall complexity of a language (Bentz et al., 2016, 2017; Baerman et al., 2015, 2017), and such measures may be under considerable debate (see Bentz et al., 2023). Additional research suggests that word complexity can be measured by word length (see Lewis & Frank, 2016; and see Piantadosi et al., 2011; Meylan & Griffiths, 2021; Koplenig et al., 2022 for further discussion in relation to frequency in corpora). Words which are formed from longer sequences of phonemes are more difficult to process, and this is also related to their conceptual complexity. The fact that word length is related to a word’s complexity is important for our understanding of word order, i.e. how closely is the ordering of constituents related to how we process words/structures of varying lengths?
Research into how speakers process words has identified several constraints and tendencies that have various implications for how languages develop. Hawkins (2014) summarizes this literature in arguing for a comprehensive theory (the Performance-Grammar Correspondence Hypothesis; PGCH) based firmly in the idea of efficiency as an explanation for a large number of observed patterns and historical changes cross-linguistically. [^6] Levshina (2023) argues for another explanation termed “communicative” efficiency, which identifies “accessibility” as a factor in language use. This seems to overlap to some degree with the effects identified by “surprisal”. As Levshina’s observations do not make specific predictions about word order crosslinguistically, I revisit them in the discussion. His essential claim is that speakers minimize complexity (in forms, distance) in order to speed up processing, a claim that aligns with observations about constituent “weight” (Wasow, 2002; Zipf, 1949) and “dependency length minimization” (Futrell et al., 2015). However, this claim contrasts with findings from Surprisal Theory (Hale 2001; Levy 2008) that speakers maximize complexity, i.e. that more complex forms are preferred by speakers since they support information content/retrieval and reduce entropy in utterances (see Rezaii et al. 2023).
These two opposing perspectives on processing (the minimization vs maximization of complexity) make two different predictions for word order crosslinguistically. The PGCH claim (and Wasow’s) is that “heavier” or more complex constituents are typically “put off” until later in the sentence, while less complex constituents are put first. This predicts that a language in which nouns in general are easier to process than verbs should display SV order, while one in which verbs are easier to process should display VS order. Yet evidence from languages like Japanese, which places heavier phrases closer to the beginning of the sentence (Hawkins, 1994; Yamashita & Chang, 2001; Yamashita, 2002; and see discussion in Wasow, 2022) supports the opposing claim by Surprisal Theory that words with higher information content (or which improve predictability of following words) are more likely to occur first. Here the prediction is that a language with heavier nouns should display SV order, and one with heavier verbs should display VS order.
Such predictions, if true, should be observable in crosslinguistic corpora. An important note here is that while the processing literature is largely grounded in a functional/usage-based perspective, it does not necessarily require a usage-based stance. That is, cognitive processes can be argued to apply at various levels of processing, including above the “innate” bedrock of current Minimalist theory (see Rastelli 2025). If there is a universal processing bias in the brain related to the complexity of structures or words in language, as has been claimed, such a bias would apply regardless of whether grammatical rules are innate or learned, and would impact or drive how such rules are realized or change over time.
2.1 Main hypotheses
@@
Using word length as a measure of word complexity, the current paper compares average lengths of particular classes of words (nouns/arguments and verbs/predicates) with word order in individual languages. The two processing claims in the literature give rise to the following two competing hypotheses regarding this relationship:
@@
2.2 Investigating word order and length
Using the 1,597 languages in the tagged PBC, we can identify basic word order via the N1 ratio. [^7] As an example, English [ISO 639-3: eng] has an N1 ratio of 4.51 and is classed as SV, whereas Irish [ISO 639-3: gle] has an N1 ratio of 1.24 and is classed as VS. Details on the N1 ratio and its validity are given above and at the linked repository (Fn 4). We can also identify the average length of nouns/arguments and verbs/predicates in each language’s corpus and see how well this correlates with the N1 ratio.
@@ 2.2.1 Word order
Of the languages in the tagged PBC, 976 are also found in typological databases with categorizations for word order. For the remainder, the work described at the linked repository has shown that the N1 ratio is a valid corpus-based proxy for word order. Accordingly I used the N1 ratio to impute the (SV, VS, free) category of the remaining 621 languages in the tagged PBC not categorized by WALS, AUTOTYP or Grambank. [^8] I trained a Gaussian Naive Bayes classifier on the known data to predict the category (SV, VS, free) of languages for which word order status was unknown, based on the N1 ratio as derived from the PBC. Code to reproduce these results can be found at the linked repository (Fn 4). This resulted in 1,251 languages classified as “SV”, 295 languages classified as “VS”, and 51 languages classified as “free”.
@@ 2.2.2 Nouns and verbs
To measure word length, I counted the number of characters in a word tagged as Noun/Argument or Verb/Predicate and computed an average length for each class on a per-language basis. [^9] For example, English [ISO 639-3: eng] arguments have an average length of 5.9 characters, while English predicates have an average length of 5.75. Irish [ISO 639-3: gle] arguments have an average length of 6.31 characters, while Irish predicates have an average length of 6.29 characters. In the tagged PBC corpus the average number of (unique) arguments per language is 732, while the average number of predicates is 495. The mean length of arguments is 6.76 characters, while the mean length of predicates is 6.90 characters. Conducting a Student’s T-Test showed that this is statistically significant (p = 0.006), indicating that crosslinguistically, arguments are shorter than predicates. All statistics showed a bimodal (normal) distribution.

### v2

@@ Abstract
A fundamental concern in linguistics has been to understand how languages change, such as in relation to word order. Since the order of words in a sentence (i.e. the relative placement of Subject, Object, and Verb) is readily identifiable in most languages, this has been a productive field of study for decades (see Greenberg 1963; Dryer 2007; Hawkins 2014). However, a language’s word order can change over time, with competing explanations for such changes (Carnie & Guilfoyle 2000; Crisma & Longobardi 2009; Martins & Cardoso 2018; Dunn et al. 2011; Jäger & Wahle 2021). This paper proposes a general universal explanation for word order change based on a theory of communicative interaction (the Min-Max theory of language behavior) in which agents seek to minimize effort while maximizing information. Such an account unifies opposing findings from language processing (Piantadosi et al. 2011; Wasow 2022; Levy 2008) that make different predictions about how word order should be realized crosslinguistically. The marriage of both “efficiency” and “surprisal” approaches under the Min-Max theory is justified with evidence from a massive dataset of 1,942 language corpora tagged for parts of speech (Ring 2025), in which average lengths of particular word classes correlates with word order, allowing for prediction of basic word order from diverse corpora. The general universal pressure of word class length in corpora is shown to give a stronger explanation for word order realization than either genealogical or areal factors, highlighting the importance of language corpora for investigating such questions.
@@ 1 Introduction
There are two major approaches to studying word order. The first (typological) approach is to observe word orders in natural languages and identify patterns whereby the orders of certain constituents are correlated with other grammatical patterns. This approach has developed multiple crosslinguistic databases with expert determinations of word order (WALS, Dryer & Haspelmath 2013; AUTOTYP, Bickel et al. 2023; Grambank, Skirgård et al. 2023) and has identified statistical universals in which the orders of certain elements increase the likelihood for other patterns to be present in a language (Greenberg 1963; Dryer 2007, 1992, 2011; Verkerk et al. 2025). The second (processing) approach focuses on experimental evidence for how constituents in particular languages are processed in the brain. This approach has identified potential universal tendencies or preferences in processing, notably a preference for “agents first” (Culbertson et al. 2012; Futrell et al. 2015a; Sauppe et al. 2023), “efficiency” (Zipf, 1949; Wasow, 2002; Futrell et al., 2015b), and “surprisal” (Hale 2001; Levy 2008).
While the typological observations do not make predictions about a language’s basic word order (the universals take basic word orders as a starting point and make further predictions, see Verkerk et al. 2025), the processing research does imply some particular claims and predictions about how words should be ordered in languages of the world. For example, the “agents first” (or “Subject first”) bias found in experiments suggests that there is a universal psychological pressure for speakers of all languages to produce agent arguments (“do-ers” of actions, typically represented by animate nouns and pronouns) before other kinds of constituents in sentences. This aligns with observations of the dominant crosslinguistic word order being Verb-final or Subject-initial (Schouwstra & de Swart 2014; Schouwstra et al. 2022).
At the same time, findings from language processing studies suggest that other pressures mediate the realization of word order patterns, namely in relation to “efficiency” and “surprisal”. The efficiency literature largely follows from Zipf’s (1932; 1949) observations that a) languages show a logarithmic distribution whereby the most frequently used words occur much more often than the next most frequent, and that b) the most frequent words in a language are also the shortest. The observation that shorter (or less complex) elements are more likely to occur initially in a sentence, while longer (or more complex) elements occur later, can be traced to Behaghel (1909), and the general idea is that this supports efficient communication. In contrast to this view, the surprisal literature focuses on how salient or predictable items are, suggesting that longer (or more complex) words are more likely to occur first in a sentence since they improve predictability of following constituents (Hale 2001; Levy 2008; Sauppe et al. 2021). This is supported by clinical evidence (Rezaii et al. 2023) and information theoretic studies (Zaslavsky 2020; Tucker et al. 2025).
Such studies have implications for particular languages regarding the ordering of constituents. For example, if a language has longer verbs than nouns, we would expect the language to be Verb-final based on the efficiency literature (longer or more complex words last), whereas the surprisal literature would predict it to be Verb-initial (longer or more complex words first). These competing predictions represent hypotheses that have yet to be tested to any great degree (though some attempts have been made, at least in relation to dependency structures, see Jing et al. 2021; Hahn & Xu 2022).
The lack of testing highlights the complexity of investigating pressures that influence the evolution and realization of language structures, as well as the dearth of datasets and databases that could address these questions. [^1] Datasets that have been used by previous research include the Universal Dependencies Treebanks (Zeman et al. 2024) and ODIN (Xia et al. 2016) projects. The UDT contains 160+ hand-annotated languages available for crosslinguistic investigation, but these are not parallel texts and the number of sentences in each language is variable, with some having fewer than 200 sentences in the v2.14 release. The ODIN project (and its re-packaged form via LECS: https://huggingface.co/datasets/lecslab/glosslm-corpus) has interlinear glossed text (IGT) from over 1,800 languages, but the majority of these are single-word/line examples, and only 300 of the languages in their dataset have more than 100 sentences. Additional cross-linguistic investigations using tagged corpora include Hahn et al. (2020: 51 languages), Jing et al. (2021: 71 languages), and Hahn et al. (2022: 80 languages). For each of these studies the majority of languages are from the Indo-European language family. Besides limiting generalizability of findings, such data risks bias toward a particular group of languages (Blasi et al. 2022), in much the same way that other scientific findings have historically been biased toward WEIRD demographics (Henrich et al. 2010). In particular, questions regarding the length of constituents in relation to their order in sentences requires a large number of sentences tagged for parts of speech (at minimum). Given that languages develop over time via cultural evolution, it is also difficult to separate general cognitive pressures from historical (genealogical) or areal factors. This means that inferring crosslinguistic or universal principles necessitates both a large number of languages as well as sufficient diversity of languages to establish statistical validity.
Additionally, the fact that languages develop over time through communicative interaction requires that any explanation for universal principles be related to such interactions. That is, in an evolutionary linguistic paradigm it is individuals (speakers) in a population (language) that drive variation and select particular linguistic variants for propagation. The selection process itself is likely motivated by various factors, two of which are combined in the Min-Max theory of language behavior. The Min-Max theory proposes that in production speakers select variants based on two competing pressures: the pressure to minimize effort and the pressure to maximize information (Figure 1). This language behavior is considered to be core to what drives selection and propagation of linguistic constructions, and thus change over time (in, i.e. word order).
Figure 1: Min-Max schema
The current paper takes advantage of a massive tagged parallel corpus to show how these pressures are realized in actual language data in relation to word order. While the dataset is still under development, the baseline data is tagged for parts of speech, which allows us to derive a proxy measure for word order from individual language corpora, as well as to measure other linguistic properties such as the average length of various word types. Results show that basic word order in intransitive clauses is highly correlated with the length of particular word classes, allowing for the prediction of a language’s word order based on corpora not in the dataset and explaining variance better than both genealogy and areal factors.
In the remainder of this paper I first highlight the properties of the dataset (§2). I then discuss three studies that investigate word length in relation to word order (§2.1), the predictive validity of word lengths for word order (§2.2), and the variance predicted by word length in regression models (§2.3). I conclude with a discussion of the implications of these studies in relation to language evolution (§3), presenting the Min-Max theory of language behavior (§3.1) as well as some limitations and future directions regarding the current effort (§3.2).
2 Methodology and hypotheses
The taggedPBC (Ring 2025) [^2] https://github.com/lingdoc/taggedPBC - see fn 5 for a link to a minimal OSF repository that allows for replication of the studies in this paper. contains pos-tagged parallel text data in 1,942 languages (roughly 25% of the world’s languages), with over 1,800 parallel verses for more than 1,850 of those languages. The dataset represents 155 language families and 78 isolates, dwarfing previously available resources, and the tagging methodology (via crosslingual tag transfer) has been shown to correspond well with existing automated taggers for high-resource languages (as found in the SpaCy and TranKit NLP libraries). [^3] https://spacy.io ; https://nlp.uoregon.edu/trankit Additionally, a novel measure extracted from individual corpora in this dataset (the “N1 ratio”) is significantly correlated (p < 0.001) with expert determinations of a language’s basic word order for intransitive clauses (“SV”, “VS”, “free”; Figure 2) as identified in the typological databases WALS (Dryer & Haspelmath 2013), Grambank (Skirgård et al. 2023), and Autotyp (Bickel et al. 2023). [^4] The procedure for determing the N1 ratio is essentially to count the number of sentences in a corpus that have both arguments (nouns, pronouns, etc) and predicates (verbs, auxiliaries), and then to derive a ratio from the number of argument-first sentences divided by the number of predicate-first sentences in that set. As an example, English [ISO 639-3: eng] has an N1 ratio of 4.51, whereas Irish [ISO 639-3: gle] has an N1 ratio of 1.24, based on their respective corpus in the taggedPBC. This ratio can then be related to other aspects of language, such as word order classifications. See Ring (2025) for further details. What this means is that using this dataset we can derive word order and word length information in a systematic manner, allowing for an investigation of word length in relation to word order.
Figure 2: The N1 ratio and word order in 3 typological databases (p < 0.001; Ring 2025)
The following studies focus on intransitive word order (i.e. the basic word order of intransitive clauses, in which a predicate requires a single “S” argument) in languages found in the taggedPBC. [^5] Minimal data and Python code to replicate the results in this paper can be found at https://osf.io/nsaq2 (3MB). More comprehensive data and code (7GB) can be found at the taggedPBC github repository (fn 2). The first study measures word classes, showing that average lengths of nouns and verbs in a language’s corpus are highly correlated with SV, VS, and free word order classifications when frequency is considered. The second study investigates whether this effect occurs outside of the taggedPBC, showing that information on average lengths of nouns and verbs computed from different corpora also allow for the classification of languages as SV or VS, differentiating between historical and modern languages for which word order changes are known. The third study examines whether the lengths of nouns and verbs in corpora provides a better indication of word order than known areal or genealogical signals, and shows that in heirarchical linear regressions, word length is indeed a better predictor of word order.
2.1 Study 1: Word length and word order
@@
Study 1 follows previous studies in using word length as a measure of word complexity (Lewis & Frank 2016; Bentz et al. 2016, 2017; Baerman et al. 2015, 2017; Bentz et al. 2023; and see Piantadosi et al., 2011; Meylan & Griffiths, 2021; Koplenig et al., 2022 for further discussion in relation to frequency in corpora), comparing average lengths of particular classes of words (nouns/arguments and verbs/predicates) with word order in individual languages. The two processing claims in the literature (from “efficiency” and “surprisal” approaches) give rise to the following two competing hypotheses regarding this relationship:
@@
For the 1,942 languages in the taggedPBC, we can identify basic word order via the N1 ratio. We can also identify the average length of nouns/arguments and verbs/predicates in each language’s corpus and see how well this correlates with the N1 ratio.
@@ 2.1.1 Word order
Of the languages in the taggedPBC, 1,152 are also found in typological databases with categorizations for word order. For the remainder, the N1 ratio was used to impute the (SV, VS, free) category of the remaining 790 languages in the taggedPBC not categorized by WALS, AUTOTYP or Grambank, and the final categorizations were stored for all 1,942 languages. [^6] I trained a Gaussian Naive Bayes classifier on the known data to predict the category (SV, VS, free) of languages for which word order status was unknown, based on the N1 ratio as derived from the taggedPBC. Code to reproduce these results can be found at the linked repository (fn 2). This resulted in 1,548 languages classified as “SV”, 331 languages classified as “VS”, and 63 languages classified as “free”.
@@ 2.1.2 Nouns and verbs
To measure word length, I counted the number of characters in a word tagged as Noun/Argument or Verb/Predicate [^7] Here I conducted several tests to see if there were different effects for different word types and groupings. “Argument” is a grouping of words tagged as NOUN, PRON, PROPN in the universal tag set, while “predicate” refers to a grouping of VERB and AUX tags. and computed an average length for each class on a per-language basis. [^8] For example, English [ISO 639-3: eng] arguments have an average length of 5.9 characters, while English predicates have an average length of 5.75. English nouns have an average length of 5.9 characters, while English verbs have an average length of 5.8 characters. Irish [ISO 639-3: gle] arguments have an average length of 6.31 characters, while Irish predicates have an average length of 6.29 characters. Irish nouns have an average length of 5.2 characters, while Irish verbs have an average length of 7.9 characters. In the taggedPBC the average number of (unique) arguments per language is 741, and the average number of predicates is 498. The mean length of arguments is 6.71 characters, and the mean length of predicates is 6.82 characters. Conducting a Student’s T-Test showed that this is statistically significant (p < 0.001), indicating that crosslinguistically, arguments are shorter than predicates. All statistics showed a normal distribution.